OpenStudy (anonymous):

in a large bag of candy each of the 5 colors red orange yellow blue and green occurs with the same probabilty you reach in and select 1 candy find P (yellow or orange)

OpenStudy (p0sitr0n):

OR means you want the sum of the probabilities. thus, P(y or O)=P(Y)+P(Y)-P(Y and O). SInce those do not intersect (are disjoint or mutually exclusive) then P(Y or O) = 1/5+1/5=2/5